OpenStudy (anonymous):

Please Help. Classify the system as Independant, dependant or inconsistant. Problem in comments. best answer and will become a fan.

OpenStudy (anonymous):

\[3x+6y=15\]\[2x+4y=3\]

OpenStudy (anonymous):

lets multiply 2x+4y=3 by 1.5 we get 3x+6y=4.5 and we have 3x+6y=15 so the system is obviously inconsistent because we 0 is not equal to -10.5

OpenStudy (anonymous):

we know*

OpenStudy (anonymous):

thank you for so much help :)